(I) You have just been appointed the product manager of the "FIFO" electric blankets in a large consumer products company. As part of your new job, you want to develop an understanding of the financial situation for your product. Your brand assistant has provided you with the following facts:

a. Retail selling price                                                  $40 per unit

b. Retailer's margin                                                     25%

c. Jobber's margin                                                       12%

d. Wholesaler's margin[1]                                              20%

e. Direct factory labor                                                $2 per unit

f. Raw materials                                                         $1 per unit

g. All factory and administrative overheads              $2.7 per unit (if unit volume = 100,000)

h. Salesperson's commissions                                    10% of manufacturer's selling price

i. Sales force travel costs                                            $232,000

j. Advertising                                                             $800,000

k. Total market for electric blankets                          1 million units

l. Current yearly sales of "FIFO"                         210,000 units

Questions

1. What is the contribution per unit for the "FIFO" brand?[2]

2. What is the break-even-volume in units and in dollars?

3. What market share does the FIFO brand need to break even?

4. What is the current total contribution?

5. What is the current before-tax profit of the FIFO brand?

6. What market share must FIFO obtain to contribute a before tax profit of exactly $6 million?

[1] The wholesaler sells to the jobber who, in turn, sells to the retailer.

[2] You are a part of the company that manufactures FIFO. Hence, you have to view this problem   from the perspective of the manufacturer of the product (and not from that of the middlemen).

following facts.

a. Retail selling price                                                  $60 per unit

b. All margins the same as before

c. Direct factory labor                                                $3 per unit

d. Raw materials                                                        $5 per unit

e. Additional factory and admin. overheads              $3.5 per unit

                                                                                    (if unit volume = 50,000)

f. Salesperson's commissions: the same percent as before

g. Incremental sales force travel cost                         $68,000

h. Advertising for Super FIFO $600,000

i. New equipment needed                                          $800,000 (to be depreciated over 10 years)

j. Research and development spent                           $200,000

            up to now

k. Research and development to be                           $600,000 (to be amortized over 5 years)

     spent this year to commercialize

     the product

Questions

1. What is the contribution per unit of the Super-FIFO brand?

2. What is the break-even volume in units and in dollars?

3. What is the sales volume in units necessary for Super FIFO to yield in the first year, a 25          percent return on the equipment to be invested in the project?

(III) The $60 selling price for Super FIFO seems high to you. You thought you might lower the price to $50 per unit and raise retail margin to 30 percent.

Question

            What is the break-even volume in units?
Answer #1
Part - 1
1. Contribution p.u for FIFO Brand
Particulars Amount($ p.u)
Retail Selling Price 40
Less : Retailer's Margin 10
Jobber's Selling Price 30
Less : Jobber's margin 3.6
Wholesaler's Selling Price 26.4
Less : Margin 5.28
Manufacturer's Selling Price 21.12
Less : Variable Costs
Raw Material 1
Direct Labor 2
Factory & Admin Overhead 2.7
Commission (10% of MSP) 2.11
7.812
Contribution p.u 13.31
2. Break-even Volume
Break-even Volume = Fixed Cost / Contribution p.u
Break-even Volume = (800000+232000) / 13.31
Break-even Volume = 77535.69
Break-even Sales = 77535.69*21.12
Break-even Sales = $ 1,637554
3. Market Share
Total Market required = Break-even Volume / Total Market for counter top water dispensers
Total Market required = 77535.69 / 1000000
Total Market required = 7.75%
4. Current Total Contribution
Current Total Contribution = Contribution p.u. * Total Sales
Current Total Contribution = 13.31 * 210000
Current Total Contribution = $2795100
5. Before Tax Profit
Before Tax Profit = Total Contribution - Fixed Costs
Before Tax Profit = 2795100 - 1032000
Before Tax Profit = $1763100
6. Market Share Required
Total Contribution Required = 6000000+1032000
Total Contribution Required = $7032000
Contribution p.u. = $ 13.31
Total No. of units required = 7032000/13.31
Total No. of units required = 528324.6
Total market share = 528324.6 / 1000000
Total market share = 52.83%
Part 2
1. Contribution p.u for Super FIFO Brand
Particulars Amount($ p.u)
Retail Selling Price 60
Less : Retailer's Margin 15
Jobber's Selling Price 45
Less : Jobber's margin 5.4
Wholesaler's Selling Price 39.6
Less : Margin 7.92
Manufacturer's Selling Price 31.68
Less : Variable Costs
Raw Material 5
Direct Labor 3
Factory & Admin Overhead 3.5
Commission (10% of MSP) 3.17
14.67
Contribution p.u 17.01
2. Break-even Volume
Break-even Volume = Fixed Cost / Contribution p.u
Break-even Volume = (300000+600000+800000/10+200000+600000/5) / 17.01
Break-even Volume = 76425.63
Break-even Sales = 76425.63* 31.68
Break-even Sales = $ 2421164
3. Sales volume required
Expected return = 800000*25%
Expected return = 200000
Sales Volume = (Expected return + Fixed Cost)/Contribution p.u
Sales Volume = (200000+1300000)/17.01
Sales Volume =88183.42
Part - 3
Particulars Amount($ p.u)
Retail Selling Price 50
Less : Retailer's Margin 15
Jobber's Selling Price 35
Less : Jobber's margin 4.2
Wholesaler's Selling Price 30.8
Less : Margin 6.16
Manufacturer's Selling Price 24.64
Less : Variable Costs
Raw Material 5
Direct Labor 3
Factory & Admin Overhead 3.5
Commission (10% of MSP) 2.46
13.96
Contribution p.u 10.68
Break-even Volume = Fixed Cost / Contribution p.u
Break-even Volume = (300000+600000+800000/10+200000+600000/5) / 10.68
Break-even Volume = 140449.4